a) Evaluate the following two feasible investments A and B having different useful life, if MARR is
15 % per year. Use PW method with repeatability assumptions. 8

Investment of A (Rs.) Investment of B (Rs.)

Investment 40,000 50,000
Net annual
revenue

15,000 20,000
Salvage value 5,000 6,000
Useful life 3 years 5 years
1
Expert's answer
2020-05-27T09:21:28-0400

"PW_B(A)=15,000(P/A,15\%,3)+5,000(P/F,15\%,3)"

"=15,000×\frac{(1+0.15)^3-1}{0.15×(1+0.15)^3}+5,000×(1+0.15)^3=37,536"

"PW_C(A)=40,000"

"B/C(A)=\frac{PW_B(A)}{PW_C(A)}=\frac{37,536}{40,000}=0.94"

"PW_B(B)=20,000(P/A,15\%,5)+6,000(P/F,15\%,5)"

"=20,000×\frac{(1+0.15)^5-1}{0.15×(1+0.15)^5}+6,000×(1+0.15)^5=70,026"

"PW_C(B)=50,000"

"B/C(B)=\frac{PW_B(B)}{PW_C(B)}=\frac{70,026}{50,000}=1.40"